KaylaCortez:

If the sin 0° = 0, then which statement is true? a. cos 90 = 1 because the cosine and sine are complements b. cos 180 = 1 because the cosine and sine are supplements c. cos 90 = 0 because the cosine and sine are complements d. cos 180 = 0 because the cosine and sine are supplements

justjm:

Apply the complement rule here \[sin(\theta)=cos(\frac{\pi}{2}-\theta)\]

justjm:

It is given to you here that theta = 0, so just plug that in and that should direct you to the correct choice